# On $t$-structures adjacent and orthogonal to weight structures

### v1

@@ 4.2 More on the construction of $t$-structures
Then there exists a $t$ structure $t$ (left) adjacent to $w$ (on $\underline{C}$). Its heart is equivalent to the category of $R$-linear functors from ${\underline{Hw}}^{op}$ into the category of finitely generated $R$-modules.
@@
3. The first part of the assertion is a particular case of [Nee18, Corollary 0.5]; note that in this case $D^{perf}(X)$ is equivalent to the bounded derived category of $X$; cf. Remark 1.4(3) of [Bon24].
@@
1. In all the examples of $R$-saturated categories known to the author one can achieve the same result by means of applying (the somewhat more clumsy) Corollary 4.2.5 instead; see Theorem 4.2.2(1) and Proposition 4.2.6 of [Bon24].
2. Similarly to Remark 3.2.4(3) one can apply Corollary 4.2.2(1,2) to semi-orthogonal decompositions. One obtains that if $\underline{C}$ is saturated, $\underline{C}^{\prime}=\underline{C}_{\underline{A}^{T}}$ (note that $\underline{C}^{\prime}=\underline{C}$ if $T=\operatorname{Spec}R$), and $(\mathcal{A},\mathcal{B})$ is a semi-orthogonal decomposition of $\underline{C}$, then the couple $((\operatorname{Obj}\mathcal{A})^{\perp}\cap\operatorname{Obj}\underline{C}^{\prime},(\operatorname{Obj}\mathcal{B})^{\perp}\cap\operatorname{Obj}\underline{C}^{\prime}=\mathcal{A}\cap\operatorname{Obj}\underline{C}^{\prime})$ is a a semi-orthogonal decomposition of $\underline{C}^{\prime}$. Consequently, if $\mathcal{A}$ is a left admissible subcategory of $\underline{C}$ then it is also admissible in it. This statement generalizes Proposition 2.6 of [BoK89], where the case $R=\mathbb{C}$ was considered.
Now, it is well-known that the categories of the type $D_{perf}(X)$ often admit non-trivial semi-orthogonal decompositions (in particular, in the case where $R$ is a field). Hence there can exist unbounded weight structures on them; see Remark 4.1.5(2). On the other hand the author suspects that any weight structure on $D_{perf}(X)$ is essentially bounded; hence one can apply Corollary 4.2.2(1,2) to it.
@@
3. Clearly, if $\underline{C}$ is $R$-saturated and the category $\underline{A}$ is a weak Serre subcategory of $R-\operatorname{Mod}$ then the corresponding category $\underline{C}_{\underline{A}}$ depends on $\underline{A}\cap R-\operatorname{mod}$ only. Now, $\underline{A}\cap R-\operatorname{mod}$ is a weak Serre subcategory of $R-\operatorname{mod}$ (cf. Remark 1.1), and any weak Serre subcategory of $R-\operatorname{mod}$ of $R-\operatorname{mod}$ consists of finitely generated $R$-modules supported at some $T$ as in Corollary 4.2.2(2); see Theorem A of [Tak08]. [^7] Note that extension closed abelian subcategories are called coherent ones in (Definition 2.3(1) of) [Tak08]. Consequently, it does not make sense to consider any $\underline{A}$ distinct from $\underline{A}^{T}$ in Corollary 4.2.2(2).
@@
Assume that a scheme $X$ is projective (see Proposition 3.3.5) over $\operatorname{Spec}R$ (where $R$ is Noetherian); take $\underline{C}=D(\operatorname{Qcoh}(X))$ (the derived categories of quasicoherent sheaves on $X$), and ${\underline{C}^{0}}=\underline{C}^{({\aleph_{0}})}=D_{perf}(X)$ (see Proposition 3.3.5 once again). 1. Then the intersection subcategories $\underline{C}_{{\aleph_{0}}}\cap\underline{C}^{{}^{\prime}0}_{+}$, $\underline{C}_{{\aleph_{0}}}\cap\underline{C}^{{}^{\prime}0}_{-}$, and $\underline{C}_{{\aleph_{0}}}\cap\underline{C}^{{}^{\prime}0}_{b}$ equal $D^{+}_{coh}(\operatorname{Qcoh}(X))$, $D_{coh}^{-}(\operatorname{Qcoh}(X))$, and $D_{coh}^{b}(\operatorname{Qcoh}(X))$, respectively; here a complex $N\in D_{coh}(\operatorname{Qcoh}(X))$ belongs to $D^{+}_{coh}(\operatorname{Qcoh}(X))$ (resp. $D_{coh}^{-}(\operatorname{Qcoh}(X))$, $D_{coh}^{b}(\operatorname{Qcoh}(X))$) whenever its (coherent) cohomology sheaves $H^{i}(N)$ vanish for $i\ll 0$ and (resp. for $i\gg 0$, in both cases). 2. Take $T$ to be a union of closed subsets of $S$. Then to characterize the intersections of corresponding category $\underline{C}_{\underline{A}^{T}\cap R-\operatorname{mod}}$ (see Definition 2.3.4(2)) with $\underline{C}^{{}^{\prime}0}_{+}$, $\underline{C}^{{}^{\prime}0}_{-}$, and $\underline{C}^{{}^{\prime}0}_{b}$ one should add to the aforementioned conditions the assumption that the cohomology sheaves of $N$ are supported on $T$.
@@
1. The calculation of these categories is given by Theorem 4.2.2(3) of [Bon24].
@@
1. Thus one can apply Corollary 4.2.5 to all the intersection subcategories mentioned in Proposition 4.2.6. 2. Corollary 4.2.5 also implies that for any essentially bounded weight structure on $D_{coh}^{b}(\operatorname{Qcoh}(X))\cong D^{b}(\operatorname{coh}(X))$ (see [stacks, Tag 0FDA]) there exists an anti-orthogonal $t$-structure on $D_{perf}(X)$; see Proposition 4.2.6 of [Bon24]. The author conjectures that there do not exist bounded weight structures on $D^{b}(\operatorname{coh}(X))$ if $X$ is singular. On the other hand, one may "start from" a bounded weight structure on a component of a semi-orthogonal decomposition of $X$ (of arbitrary length; see Definition 2.3.1(II.2) of ibid.) to obtain an essentially bounded weight structure on the whole $\underline{C}$; see Remark 4.1.5(2) and Proposition 2.3.2 of ibid. 3. Applying Remark 4.1.5(2) one can prove that Corollary 4.2.5 allows to generalize all the statements in Theorem 4.2.7 of ibid. except the bijection statement in part II.1 of loc. cit. from semi-orthogonal decompositions to essentially bounded (above, below, or both) weight structures. Probably, this statement and examples for it will be studied in a succeeding paper.
@@ 5 The existence of orthogonal weight structures
In §5.2 we prove that the existence of an adjacent weight structure is closely related to the existence of enough projectives in the heart of $t$. This assumption ensures the existence of weight structure on certain ("rather large") subcategories of $\underline{C}$ that are orthogonal to $t$. Under certain assumptions, we also obtain a weight structure adjacent to $t$.
@@ 5.1 On hearts and $\pm$-orthogonality of structures
We will say that $\underline{H}$ is silting if it is connective and densely generates $\underline{C}$ (cf. §3.1 of [KoY14]).
@@
2.
$\underline{C}(P,-)\cong\underline{C}(P,H_{0}^{t}(-))\cong{\underline{Ht}}(H_{0}^{t}(P),H_{0}^{t}(-));$
the first of these transformations comes from the transformations $\operatorname{id}_{\underline{C}}\to t_{\leq 0}$ and $H_{0}^{t}\to t_{\leq 0}$ (see Proposition 1.1.5(1,3)). Moreover, $H_{0}^{t}(P)$ is projective in ${\underline{Ht}}$.
@@
I. If $0\to A_{1}\to A_{2}\to A_{3}\to 0$ is a short exact sequence in ${\underline{Ht}}$ then $A_{1}\to A_{2}\to A_{3}\to A_{1}[1]$ is a distinguished triangle according to Proposition 1.1.5(2). Applying the functor $\underline{C}(P,-)$ to it and recalling the definition of $P_{t}$ we obtain an exact sequence $\{0\}=\underline{C}(P,A_{3}[-1])\to E^{P}(A_{1})\to E^{P}(A_{2})\to E^{P}(A_{3})\to E^{P}(A_{1}[1])=\{0\}$; hence $E^{P}$ is exact indeed. Moreover, the functors $\Phi(P,-)$ and $E^{P}\circ H_{0}^{t}$ are homological and annihilate both $\underline{C}_{t\leq-1}$ and $\underline{C}_{t\geq 1}$; hence they are isomorphic.
@@
3. We should prove that any projective object $P_{0}$ of ${\underline{Ht}}$lifts" to $P_{t}$.
@@
Since $w$ is orthogonal to $t$, we have $\underline{C}^{\prime}_{w\leq 0}\subset C_{1}$ and $\underline{C}^{\prime}_{w\geq 1}\subset C_{2}$. Next, $C_{1}\perp C_{2}$, and applying Proposition 1.2.4(2) we obtain $w=(C_{1},C_{2}[-1])$; see Remark 2.3.2(1). Hence $\underline{C}^{\prime}_{w=0}=C_{1}\cap C_{2}[-1]=P^{\prime}_{t}$.
@@
3. Since all of these functors are exact (see Proposition 5.1.4(1)), it suffices to verify that for any non-zero $N\in\underline{C}_{t=0}$ there exists $M\in\underline{C}^{\prime}_{w=0}$ such that $\underline{C}(M,N)\neq\{0\}$. Now, if $M\perp N$ for all $M\in\underline{C}^{\prime}_{w=0}$ then combining Proposition 2.2.2(5) with the description of $w$-pure functors provided by Theorem 2.1.2(2) of [Bon21] we obtain $\operatorname{Obj}\underline{C}\perp N$. Combining this statement with either the -- or the $+$-orthogonality of $t$ to $w$ we immediately obtain $N=0$ (i.e., a contradiction).
@@ 5.2 On the existence of orthogonal weight structures
1. Fix $M\in\underline{C}_{t=0}$ and consider its $w$-decomposition $P\stackrel{{\scriptstyle p}}{{\to}}M\to w_{\geq 1}M\to P[1]$. Since $M\in\underline{C}_{t=0}$, Proposition 5.1.5(1) implies that $P$ belongs to $\underline{C}^{\prime}_{w\geq 0}$; hence $P$ belongs to $\underline{C}^{\prime}_{w=0}$ according to Proposition 1.2.4(7)). Next, $P\in\underline{C}^{\prime}_{w\geq 0}\subset\underline{C}_{t\geq 0}$ (see Proposition 5.1.5(1) once again); hence the object $P_{0}=t_{\leq 0}P$ equals $H_{0}^{t}(P)$ (see Proposition 1.1.5(3)). Therefore $P_{0}$ is projective in ${\underline{Ht}}$ according to Proposition 5.1.4(2).
@@
Applying Proposition 5.1.5(3) we obtain that it remains to verify that the functors of the form $E^{M}$ for $M\in\underline{C}^{\prime}_{w=0}$ give a conservative family of functors ${\underline{Ht}}\to\operatorname{Ab}$. Now, for any object $N$ of ${\underline{Ht}}$ our assumptions give the existence of a projective object $P_{0}$ of ${\underline{Ht}}$ that surjects onto it. Moreover, applying Proposition 5.1.4(2) we obtain the existence of $P\in\underline{C}^{\prime}_{w=0}$ and a morphism $h$ from $P$ such that $H_{0}^{t}(h)$ is isomorphic to this surjection $P_{0}\to N$. Hence $E^{P}(N)\neq 0$ if $N$ is non-zero, and we obtain the conservativity in question (see Remark 5.1.6). ∎
@@
1. Then the full subcategory $\underline{C}_{+}$ of $\underline{C}$ whose object class equals $\cup_{i\in{\mathbb{Z}}}\underline{C}_{t\geq i}$ is triangulated, and there exists a weight structure $w_{+}$ on $\underline{C}^{\prime}$ such that $\underline{C}_{+,w_{+}\geq 0}=\underline{C}_{t\geq 0}$ and $t$ is ---orthogonal to $w_{+}$.
@@
The only non-trivial axiom check here is the existence of $w_{+}$-decompositions for all objects of $\underline{C}$. Let us verify the existence of a $w$-decomposition for any $M\in\underline{C}_{t\geq i}$ by downward induction on $i$. The statement is obvious for $i>0$ since $M\in\underline{C}_{+,w_{+}\geq 1}=\underline{C}_{t\geq 1}$ and we can take a "trivial" weight decomposition $0\to M\to M\to 0$.
@@
Now suppose that assumption b is fulfilled. Similarly to the previous proof, it suffices to verify that for the couple $w=({}^{\perp}\underline{C}_{t\geq 1},\underline{C}_{t\geq 0})$ the corresponding decompositions exist for all objects of $\underline{C}$.
@@
2. Moreover, Proposition 5.2.1 along with Theorem 5.2.2(I.3,II) can be considered as a certain complement to Theorem 3.2.3(I). So we obtain that the class of $t$-structures adjacent to smashing weight ones is "closely related" to the one of cosmashing $t$-structures such that ${\underline{Ht}}$ has enough projectives.
@@
2. Since $t$ is right non-degenerate and $\cap_{i\in{\mathbb{Z}}}\underline{C}_{t\leq i}=\cap_{i\in{\mathbb{Z}}}\underline{C}_{t\geq i}{}^{\perp}$ (see Proposition 1.1.5(4)), we obtain $\underline{\tilde{\mathcal{C}}}^{\perp_{\underline{C}}}=\{0\}$.
@@
1. Consequently, if $\underline{C}$ is well generated, satisfies the dual Brown representability property, $t$ is a right non-degenerate cosmashing $t$-structure on $\underline{C}$ and ${\underline{Ht}}$ contains enough projectives then there exists a weight structure adjacent to $t$ if and only if there exists a set that "generates" $S$ as in Proposition 5.2.4(2); see Theorem 5.2.2(II).

### v2

@@ 4.2 On $t$-structures corresponding to essentially bounded weight structures
Then there exists a $t$ structure $t$ (left) adjacent to $w$. Its heart is equivalent to the category of $R$-linear functors from ${\underline{Hw}}^{op}$ into the category of finitely generated $R$-modules.
@@
3. The first part of the assertion is a particular case of [Nee18, Corollary 0.5]; note that in this case $D^{perf}(X)$ is equivalent to the bounded derived category of $X$ (cf. Remark 1.4(3) of [Bon24b]).
@@
1. In all the examples of $R$-saturated categories known to the author one can achieve the same result by means of applying (the somewhat more clumsy) Corollary 4.2.5 instead; see Theorem 4.7(1) and Proposition 4.11 of [Bon24b].
2. Similarly to Remark 6.5(1) below one can apply Corollary 4.2.2(1,2) to semi-orthogonal decompositions. One obtains that if $\underline{C}$ is saturated, $\underline{C}^{\prime}=\underline{C}_{\underline{A}^{T}}$ (note that $\underline{C}^{\prime}=\underline{C}$ if $T=\operatorname{Spec}R$), and $(\mathcal{A},\mathcal{B})$ is a semi-orthogonal decomposition of $\underline{C}$, then the couple $((\operatorname{Obj}\mathcal{A})^{\perp}\cap\operatorname{Obj}\underline{C}^{\prime},(\operatorname{Obj}\mathcal{B})^{\perp}\cap\operatorname{Obj}\underline{C}^{\prime}=\mathcal{A}\cap\operatorname{Obj}\underline{C}^{\prime})$ is a a semi-orthogonal decomposition of $\underline{C}^{\prime}$. Consequently, if $\mathcal{A}$ is a left admissible subcategory of $\underline{C}$ then it is also admissible in it. This statement generalizes Proposition 2.6 of [BoK89], where the case $R=\mathbb{C}$ was considered.
Now, it is well-known that the categories of the type $D_{perf}(X)$ often admit non-trivial semi-orthogonal decompositions (in particular, in the case where $R$ is a field). Hence there can exist unbounded weight structures on them; see Remark 4.1.5(2). On the other hand the author conjectures that any weight structure on $D_{perf}(X)$ is essentially bounded; hence one can apply Corollary 4.2.2(1,2) to it.
@@
3. Clearly, if $\underline{C}$ is $R$-saturated and the category $\underline{A}$ is a weak Serre subcategory of $R-\operatorname{Mod}$ then the corresponding category $\underline{C}_{\underline{A}}$ depends on $\underline{A}\cap R-\operatorname{mod}$ only. Now, $\underline{A}\cap R-\operatorname{mod}$ is clearly a weak Serre subcategory of $R-\operatorname{mod}$ and any weak Serre subcategory of $R-\operatorname{mod}$ of $R-\operatorname{mod}$ consists of finitely generated $R$-modules supported at some $T$ as in Corollary 4.2.2(2); see Theorem A of [Tak08]. [^7] Note that extension-closed abelian subcategories are called coherent ones in (Definition 2.3(1) of) [Tak08]. Consequently, it does not make sense to consider any $\underline{A}$ distinct from $\underline{A}^{T}$ in Corollary 4.2.2(2).
@@
Assume that a scheme $X$ is projective (see Proposition 3.3.5) over $\operatorname{Spec}R$ (where $R$ is Noetherian); take $\underline{C}=D(\operatorname{Qcoh}(X))$ (the derived categories of quasicoherent sheaves on $X$), and ${\underline{C}^{0}}=\underline{C}^{({\aleph_{0}})}=D_{perf}(X)$ (see Proposition 3.3.5 once again).
1. Then the intersection subcategories $\underline{C}_{{\aleph_{0}}}\cap\underline{C}^{{}^{\prime}0}_{+}$, $\underline{C}_{{\aleph_{0}}}\cap\underline{C}^{{}^{\prime}0}_{-}$, and $\underline{C}_{{\aleph_{0}}}\cap\underline{C}^{{}^{\prime}0}_{b}$ equal $D^{+}_{coh}(\operatorname{Qcoh}(X))$, $D_{coh}^{-}(\operatorname{Qcoh}(X))$, and $D_{coh}^{b}(\operatorname{Qcoh}(X))$, respectively; here a complex $N\in D_{coh}(\operatorname{Qcoh}(X))$ belongs to $D^{+}_{coh}(\operatorname{Qcoh}(X))$ (resp. $D_{coh}^{-}(\operatorname{Qcoh}(X))$, $D_{coh}^{b}(\operatorname{Qcoh}(X))$) whenever its (coherent) cohomology sheaves $H^{i}(N)$ vanish for $i\ll 0$ and (resp. for $i\gg 0$, in both cases).
2. Take $T$ to be a union of closed subsets of $S$. Then to characterize the intersections of corresponding category $\underline{C}_{\underline{A}^{T}\cap R-\operatorname{mod}}$ (see Definition 2.3.4(3)) with $\underline{C}^{{}^{\prime}0}_{+}$, $\underline{C}^{{}^{\prime}0}_{-}$, and $\underline{C}^{{}^{\prime}0}_{b}$ one should add to the aforementioned conditions the assumption that the cohomology sheaves of $N$ are supported on $T$.
@@
1. The calculation of these categories is given by Theorem 4.7(3) of [Bon24b].
@@
1. Thus one can apply Corollary 4.2.5 to all the intersection subcategories mentioned in Proposition 4.2.6.
2. Corollary 4.2.5 also implies that for any essentially bounded weight structure on $D_{coh}^{b}(\operatorname{Qcoh}(X))\cong D^{b}(\operatorname{coh}(X))$ (see [stacks, Tag 0FDA]) there exists an anti-orthogonal $t$-structure on $D_{perf}(X)$; see Proposition 4.11 of [Bon24b].
The author conjectures that there do not exist bounded weight structures on $D^{b}(\operatorname{coh}(X))$ if $X$ is singular. On the other hand, one may start from a bounded weight structure on a component of a semi-orthogonal decomposition of $X$ (of arbitrary length; see Definition 2.9(II.2) of ibid.) to obtain an essentially bounded weight structure on the whole $\underline{C}$; see Remark 4.1.5(2) and Proposition 2.10 of ibid.
3. Applying Remark 4.1.5(2) one can prove that Corollary 4.2.5 allows to generalize all the statements in Theorem 4.12(I) of ibid. from semi-orthogonal decompositions to essentially bounded (above, below, or both) weight structures. It would also be very interesting to generalize part II of ibid. (somehow).
@@ 5 The existence of orthogonal weight structures
In §5.2 we prove that the existence of an adjacent weight structure is closely related to the existence of enough projectives in the heart of $t$. This assumption ensures the existence of weight structure on certain (’rather large’) subcategories of $\underline{C}$ that are orthogonal to $t$. Under certain assumptions, we also obtain a weight structure adjacent to $t$.
@@ 5.1 On hearts and $\pm$-orthogonality of structures
We will say that $\underline{H}$ is silting if it is connective and densely generates $\underline{C}$ (see §1.1 and cf. §3.1 of [KoY14]).
@@
2. $\underline{C}(P,-)\cong\underline{C}(P,H_{0}^{t}(-))\cong{\underline{Ht}}(H_{0}^{t}(P),H_{0}^{t}(-)),$ and the first of these transformations comes from the transformations $\operatorname{id}_{\underline{C}}\to t_{\leq 0}$ and $H_{0}^{t}\to t_{\leq 0}$ (see Proposition 1.1.3(1,3)). Moreover, $H_{0}^{t}(P)$ is projective in ${\underline{Ht}}$.
@@
I. If $0\to A_{1}\to A_{2}\to A_{3}\to 0$ is a short exact sequence in ${\underline{Ht}}$ then $A_{1}\to A_{2}\to A_{3}\to A_{1}[1]$ is a distinguished triangle according to Proposition 1.1.3(2). Applying the functor $\underline{C}(P,-)$ to it and recalling the definition of $P_{t}$ we obtain an exact sequence $\{0\}=\underline{C}(P,A_{3}[-1])\to E^{P}(A_{1})\to E^{P}(A_{2})\to E^{P}(A_{3})\to E^{P}(A_{1}[1])=\{0\}$; hence $E^{P}$ is exact indeed. Moreover, the functors $\ cu(P,-)$ and $E^{P}\circ H_{0}^{t}$ are homological and annihilate both $\underline{C}_{t\leq-1}$ and $\underline{C}_{t\geq 1}$; hence they are isomorphic.
@@
3. We should prove that any projective object $P_{0}$ of ${\underline{Ht}}$ lifts to $P_{t}$.
@@
Since $w$ is orthogonal to $t$, we have $\underline{C}^{\prime}_{w\leq 0}\subset C_{1}$ and $\underline{C}^{\prime}_{w\geq 1}\subset C_{2}$. Next, $C_{1}\perp C_{2}$, and applying Proposition 1.2.4(2) we obtain $w=(C_{1},C_{2}[-1])$; see Remark 2.3.2. Hence $\underline{C}^{\prime}_{w=0}=C_{1}\cap C_{2}[-1]=P^{\prime}_{t}$.
@@
3. Since all of these functors are exact (see Proposition 5.1.4(1)), it suffices to verify that for any non-zero $P\in\underline{C}_{t=0}$ there exists $M\in\underline{C}^{\prime}_{w=0}$ such that $\underline{C}(M,P)\neq\{0\}$.
Now, the functor $H^{\underline{C}^{\prime}}_{P}$ is of weight range $[0,0]$ by Proposition 2.2.2(10). Consequently, if $M\perp P$ for all $M\in\underline{C}^{\prime}_{w=0}$ then Proposition 2.2.2(5) implies $\operatorname{Obj}\underline{C}^{\prime}\perp P$. Combining this statement with either the -- or the $+$-orthogonality of $t$ to $w$ we immediately obtain $P=0$ (i.e., a contradiction).
@@ 5.2 On the existence of orthogonal weight structures
1. Fix $M\in\underline{C}_{t=0}$ and consider its $w$-decomposition
$P\stackrel{{\scriptstyle p}}{{\to}}M\to w_{\geq 1}M\to P[1].$
Since $M\in\underline{C}_{t=0}$, Proposition 5.1.5(1) implies that $P$ belongs to $\underline{C}^{\prime}_{w\geq 0}$; hence $P$ belongs to $\underline{C}^{\prime}_{w=0}$ according to Proposition 1.2.4(7)). Next, $P\in\underline{C}^{\prime}_{w\geq 0}\subset\underline{C}_{t\geq 0}$ (see Proposition 5.1.5(1) once again); hence the object $P_{0}=t_{\leq 0}P$ equals $H_{0}^{t}(P)$ (see Proposition 1.1.3(3)). Therefore $P_{0}$ is projective in ${\underline{Ht}}$ according to Proposition 5.1.4(2).
@@ 1. Fix $M\in\underline{C}_{t=0}$ and consider its $w$-decomposition
Now we apply Proposition 5.1.5(3) and obtain that it remains to verify the following: the functors of the form $E^{M}$ for $M\in\underline{C}^{\prime}_{w=0}$ give a conservative family of functors ${\underline{Ht}}\to\operatorname{Ab}$. Now, for any object $N$ of ${\underline{Ht}}$ our assumptions give the existence of a projective object $P_{0}$ of ${\underline{Ht}}$ that surjects onto it. Moreover, applying Proposition 5.1.4(2) we obtain the existence of $P\in\underline{C}^{\prime}_{w=0}$ along with a morphism $h$ from $P$ such that $H_{0}^{t}(h)$ is isomorphic to this surjection $P_{0}\to N$. Hence $E^{P}(N)\neq 0$ if $N$ is non-zero, and we obtain the conservativity in question (see Remark 5.1.6). ∎
@@
1. Then the full subcategory $\underline{C}_{+}$ of $\underline{C}$ whose object class equals $\cup_{i\in{\mathbb{Z}}}\underline{C}_{t\geq i}$ is triangulated. Moreover, there exists a weight structure $w_{+}$ on $\underline{C}_{+}$ such that $\underline{C}_{+,w_{+}\geq 0}=\underline{C}_{t\geq 0}$, and $t$ is ---orthogonal to $w_{+}$.
@@
The only non-trivial axiom check here is the existence of $w_{+}$-decompositions for all objects of $\underline{C}$. Let us verify the existence of a $w$-decomposition for any $M\in\underline{C}_{t\geq i}$ by downward induction on $i$. The statement is obvious for $i>0$ since $M\in\underline{C}_{+,w_{+}\geq 1}=\underline{C}_{t\geq 1}$ and we can take a ’trivial’ weight decomposition $0\to M\to M\to 0$.
@@
Now suppose that assumption b is fulfilled. Similarly to the previous proof, it suffices to verify that for the couple $w=({}^{\perp}\underline{C}_{t\geq 1},\underline{C}_{t\geq 0})$ the corresponding decomposition triangles exist for all objects of $\underline{C}$.
@@
2. Moreover, Proposition 5.2.1 along with Theorem 5.2.2(I.3,II) can be considered as a certain complement to Theorem 3.2.3(I). So we obtain that the class of $t$-structures adjacent to smashing weight ones is closely related to the one of cosmashing $t$-structures such that ${\underline{Ht}}$ has enough projectives.
@@
2. Since $t$ is right non-degenerate and $\cap_{i\in{\mathbb{Z}}}\underline{C}_{t\leq i}=\cap_{i\in{\mathbb{Z}}}(\underline{C}_{t\geq i}{}^{\perp})$ (see Proposition 1.1.3(4)), we obtain $\underline{\tilde{\mathcal{C}}}^{\perp_{\underline{C}}}=\{0\}$.
@@
1. Consequently, if $\underline{C}$ is well generated, satisfies the dual Brown representability property, and $t$ is a right non-degenerate cosmashing $t$-structure on $\underline{C}$ then there exists a weight structure adjacent to $t$ if and only if ${\underline{Ht}}$ contains enough projectives and there exists a set $S$ that generates $\underline{C}_{t\geq 0}$ in the sense specified in Proposition 5.2.4(2); see Proposition 5.2.1(1) and Theorem 5.2.2(II).